Zinc reacts with hydrochloric acid based on the following equation:


If 150 g of Zn reacts with 73 g of HCl:

a. Which reactant is the limiting reagent?
b. What mass of ZnCl2 would be produced?
1
Expert's answer
2013-01-17T10:43:05-0500

Zinc reacts with hydrochloric acid based on the following equation

If 150 g of Zn reacts with 73 g of HCl:

a. Which reactant is the limiting reagent?

b. What mass of ZnCl2 would be produced?

Zn + 2HCl = ZnCl2 + H2

n(Zn) = m/M

n(Zn) = 150g/65g/mol = 2,3 mol

n(HCl) = 73g/35,5 g/mol = 2 mol

the limiting reagent in this reaction - HCl, because it isin excess

n(Zn) react = 1 mol

n(ZnCl2) = 1 mol

m(ZnCl2) = M*n

M(ZnCl2) = 65 + 35,5*2 = 136 g/mol

m(ZnCl2) = 136 g/mol * 1 mol = 136 g